OpenStudy (hoothoot2000):

indicate the equation of the line through (2, -4) and having a slope of 3/5

OpenStudy (daniellelovee):

you can use point slope formula

OpenStudy (daniellelovee):

y=mx+b

OpenStudy (daniellelovee):

y=y value x=x value m=slope b=y intercept

OpenStudy (hoothoot2000):

okay, thanks :)

OpenStudy (daniellelovee):

no problem :) if you need more help just tag me

OpenStudy (hoothoot2000):

thanks

OpenStudy (daniellelovee):

np

OpenStudy (skullpatrol):

y = mx + b is called the slope-intercept form because it involves b which is called the y-intercept While $$y - y_0 = m(x-x_0)$$ is called the point slope form